OpenStudy (anonymous):

Select the descriptions that are true of the equation y - 6 = -3 (x+2). Select all that apply. slope = 1 slope = -3 (2, -2) is a point on the line. (-2, 6) is a point on the line.

OpenStudy (the_fizicx99):

y-y1=m(x-x1) Where y1 = the first y coordinate M = Slope and x1 = the first x coordinate. Look at the equation, can you conclude?

OpenStudy (anonymous):

@tHe_FiZiCx99 It's B and D.

OpenStudy (the_fizicx99):

Yes its B and D

OpenStudy (the_fizicx99):

I just read the part that said check all that apply.